# CAAP: Context-Aware Action Planning Prompting to Solve Computer Tasks with Front-End UI Only

### v2

@@ 1 Introduction
In this paper, we present an agent model that exclusively uses visual information from the screen as input and keyboard/mouse actions as output. This approach is intuitive, given that many computer-based tasks are designed specifically for human operators. Our proposed model features a modular architecture that distinctly separates the vision model, responsible for perceiving the environment, from the decision-making model, which handles action planning. This clear separation not only enhances the agent’s functionality but also improves its adaptability across various applications. To further support and maximize the effectiveness of our agent design, we have developed a novel prompting technique called Context-Aware Action Planning (CAAP). This technique systematically organizes the contextual information required for action decisions and identifies syntactic structures that trigger the most effective Chain of Thought (CoT) (Wei et al., 2022). Our agent demonstrates superior task-solving performance on the MiniWoB++ and WebShop benchmark. It surpasses previous studies by utilizing only a minimal amount of human demonstration data, thereby illustrating its efficiency and effectiveness in handling complex computer tasks with limited training cost.
@@ 2 Related work
Humphreys et al. (2022) introduced CC-Net, a compound model of ResNet, Transformer, and LSTM that uses image and DOM data as inputs and produces policies for keyboard-mouse actions. It was trained via Supervised Learning (SL) and Reinforcement Learning (RL), covering the largest number of MiniWoB++ tasks to date (104 tasks). Furuta et al. (2023) proposed WebGUM, which combines Vision Transformer (ViT) and T5 transformers. It was trained using SL to solve both MiniWoB++ and WebShop tasks, utilizing both image and HTML inputs. Shaw et al. (2023) introduced Pix2Act, the first agent to completely eliminate reliance on DOM or HTML for solving MiniWoB++ and WebShop tasks. The Pix2Act model is built upon the Pix2Struct architecture (Lee et al., 2023) and uses only image inputs. While these studies provide valuable insights and serve as important benchmarks, their methods are impractical for real-world applications as they require around a million or more human demonstration data samples, demanding thousands of hours of human labor for training. Cheng et al. (2024) introduced SeeClick, a method based on a Large Multimodal Model (LMM). Aided by the strong reasoning capabilities of the LMM, SeeClick required a much smaller dataset of 2.8K human demonstrations to train for MiniWoB++ tasks. It fell short in overall performance, however, achieving less than a 70% success rate on the 55 tasks it covered. A common feature of all these approaches is the integration of image interpretation and action planning models within a single architecture. This rigid design poses a significant limitation: when new UI elements need to be recognized or additional actions supported for a new task, the entire architecture must be retrained.
@@ 3 Modularized GUI agent for computer tasks
Figure 1: Architecture of the front-end UI agent and its task-solving workflow.
Our agent consists of three distinct modules: a visual observer that processes and interprets image inputs, a decision-maker responsible for action planning, and an action executor that carries out the planned actions. The agent operates by repeatedly executing a cycle consisting of a series of sequential processes involving these three modules. Each of these modules is further divided into functional submodules, as depicted in Figure 1. This approach offers several advantages, including modular updates, better scalability, and parallelized processing:
Modular updates
The modular architecture provides the flexibility to perform isolated updates, minimizing disruption to the overall system and allowing for more efficient model improvements. In contrast, end-to-end monolithic agent designs, where perception and decision-making are tightly integrated within one LMM, require updates to the entire model, which is time-consuming and resource-intensive.
Scalability
The agent’s modular structure supports horizontal scalability by allowing the addition of new functional submodules to enhance overall performance. For instance, a reflection submodule (Shinn et al., 2024; Zhao et al., 2024) could be added to the decision-maker module, enabling the agent to analyze and learn from previous actions or experience. The visual observer module could be supplemented with multiple vision models, each specialized for different domains or tasks. This allows the system to dynamically select and apply the appropriate model based on the context, significantly enhancing the agent’s adaptability to various environments.
Parallelization
The agent’s modular design also supports parallel processing of image models. For example, the screen image can be divided into multiple sections, with multiple instances of visual observers processing different regions concurrently. This parallel processing can be integrated seamlessly without disrupting the performance of the decision-making modules.
@@
The visual observer module employs a two-step approach to transform visual information from a screen into the linguistic domain. Initially, in the detection stage, UI elements are identified, and their spatial locations are extracted. Subsequently, these elements undergo further analysis through advanced object understanding techniques, during which detailed attributes of each element are captured. This processing results in the information being organized into structured text, ready to be handed to the decision-maker as the representation of the current state.
**footnotetext: Notes on the potential use of an LMM: We anticipate that in the near future, LMMs may advance enough to replace our current visual observer. The adoption of LMMs would reduce the need for training vision models tailored to specific domains, further simplifying the application of our agent design across various practical use cases. However, our CAAP prompting technique, which separates sensing and reasoning into distinct modules, will continue to be superior to end-to-end systems where a single LMM manages both observation and decision-making tasks. This modular approach will maintain significant advantages, including easier updates, improved scalability, and enhanced parallelization.
@@
(a) A MiniWoB++ screenshot masked in a Pix2Struct style, hiding the target element with an ‘X’ box.
(b) A semi-masked screenshot around the target element to be processed by the Visual Observer.
(c) Extracted features of the target element.
Figure 2: The masking methods used by the CAAP agent.
@@
We use the Pix2Struct model (Lee et al., 2023) to extract detailed attributes from all UI elements identified during the detection stage. We fine-tune the original Pix2Struct model to match the specific output format required for our task. During this fine-tuning process, we employ a different masking technique from Pix2Struct style, as depicted in Figure 2 (a). Our masking technique highlights the target UI element by outlining its region and darkening the surrounding image areas, as demonstrated in Figure 2 (b). This method preserves the pixel-level details of the targeted element while helping the model learn the spatial relationships between the element and its adjacent elements. The structured output from the model is presented as shown in Figure 2 (c), with the range of extractable attributes for each element detailed in Table 7 of Appendix C.
3.2 Decision maker
Upon receiving the representation of the current screen, the decision maker module is employed to determine possible actions for completing the task. The decision-making process is facilitated by its two key components: a context collector and a CAAP prompter. The context collector gathers vital contextual information that aids in task resolution. This includes exemplary demonstrations, the ultimate task goal, the history of actions performed so far, and the range of possible actions that the language model can choose from, all formatted as text. The CAAP prompter then uses this information to construct prompts that are sent to the language model, eliciting responses that guide the next steps. The CAAP prompting method, one of our major contributions in this paper, is discussed in detail in Section 4.
@@ 3.3 Action executor
The actions determined by the decision maker are implemented through a low-level interface to effect changes in the environment by the last module, action executor. It first identifies a series of recommended actions by parsing the response text from the decision maker. Then, it executes these actions using the keyboard and mouse interfaces of the operating system. After each action, it receives an updated screenshot, halting further actions if a change in the environment is detected. The types of actions that can be executed include mouse actions such as clicking, pointing, and scrolling, as well as keyboard actions like typing and using common shortcuts. These are enumerated in Table 8 of Appendix D and implemented as Python functions using the PyAutoGUI package.
@@ 4 CAAP prompting
Figure 3: Content design for the CAAP prompting. Figure 8 and 9 in Appendix H show an example of a prompt generated during the MiniWob++ experiments and the LLM’s response to this prompt.
The performance of LLM-based agents is heavily dependent on how the prompts are structured. By focusing on delivering relevant contextual information in the most efficient manner and inducing CoT reasoning during response generation, our prompting technique significantly enhances the decision-making capabilities of LLM agents in computer tasks. CAAP prompting constructs the prompt using the content depicted in Figure 3.
@@
Providing few-shot exemplars within the prompt to an LLM is a common practice that enhances the advanced reasoning capabilities of LLMs. Our CAAP prompt begins with few-shot examples, which are sequences of actions performed by human demonstrators solving tasks similar to the given objective. In practical scenarios, the Retrieval-Augmented Generation (RAG) method can be employed to extract similar task records from a database. The effectiveness of this demonstration portion of the prompt can be significantly enhanced by incorporating rationale texts behind each action, even when those rationales are not provided by the demonstrators themselves but are autonomously generated by an LLM, as shown in Section 5.3.
@@
This second part of the CAAP prompt forms the main section, containing all the essential information crucial for the language model to make accurate decisions. It includes a description of the target task, a record of past actions, state descriptions from the visual observer, and the available range of actions for the agent.
@@
The CoT-inducing instructions section contains phrases that encourage the LLM agent to interpret the surrounding context information independently, thereby inducing the Chain of Thought. Instruction for reviewing action trajectory encourages the agent to review its actions taken so far, enhancing its understanding of the connection between these past actions and the current state of the environment. Instruction for reviewing visual observation prompts the agent to examine the current state’s observations, directing its focus on relevant objects needed to complete the task. Instruction for improving action plan requests the development of a follow-up action plan, considering the self-generated texts so far. Lastly, the Instruction for deciding next actions provides guidance on the response format and requests a list of actions to be taken in the next steps.
@@ 5.1 MiniWoB++
MiniWoB++ is an extension of the OpenAI MiniWoB benchmark, introduced in 2017, designed to evaluate an agent’s ability to handle a wide range of scenarios that may occur on a computer screen (Liu et al., 2018; Shi et al., 2017; Min, ). It consists of over a hundred of web interaction tasks that challenge the agent to devise and execute effective sequences of actions. These tasks include various UI interactions, such as button clicks, form filling, and drag-and-drop operations. We utilize the MiniWoB++ benchmark to thoroughly assess our agent’s effectiveness in navigating and solving diverse tasks within a simulated computer environment.
Previous studies using the MiniWoB++ benchmark often choose subsets of tasks most suitable for demonstrating their methods. We have also selected a total of 73 MiniWoB++ tasks that meet the following criteria: (1) tasks that consist of UI elements of types recognizable by our visual observer module, as listed in Table 7 of Appendix C (2) tasks where none of UI elements extend beyond the default screen area of 160x210 pixels (3) tasks that do not require color recognition, as our visual observer module is not yet trained for this capability.
@@
For the decision-making language model, we utilize the Azure OpenAI gpt-4-0125 model. The Azure OpenAI API features the ‘function calling’ capability, thus our prompt does not directly include the list of action types (as depicted in Figure 3). Instead, this contextual information is conveyed to the language model through arguments in the function calling feature. An example of how the function calling is applied is illustrated in Figure 10 in Appendix I.
@@
Table 1 compares the performance of our approach with other studies reported on the MiniWoB++ benchmark. Since each study reports results over different sets of tasks (See Figure 11 of Appendix K.), we have included both the SRs reported in the literature and the number of tasks used in each case. First, CC-Net, which covers the largest set of tasks, was trained on 2.4M human demonstration data across 104 tasks. However, when using only image input (excluding DOM), its performance dropped to 24%, indirectly highlighting the challenge of solving computer tasks using image inputs alone. Pix2Act, trained on 1.3M human demonstrations across 59 tasks, achieved an SR of 96.2%, demonstrating the feasibility of agents solving computer tasks with image input only. SeeClick, despite using a relatively small dataset of 2.8K demonstrations to learn click locations on screenshots, achieved an SR of only 69.4% across 55 tasks, showing limited effectiveness. In contrast, our approach, using only 97 demonstration samples, achieved a high task SR of 94.5% across 73 tasks. Although an additional 1.76K samples***These image samples were annotated via a partially automated process, making the data preparation process relatively efficient. were required to train the model for image input interpretation due to the modular nature of our approach, our method still demonstrated its effectiveness on a broader range of computer tasks with about 1.86K samples in total. Task-level SRs are detailed in Table 12 of Appendix J.3.
Of the 73 tasks included in our experiments, six tasks—drag-circle, drag-single-shape, find-greatest, generate-number, odd-or-even, and sign-agreement—were excluded from agent’s training, as they were not covered in the other studies. These tasks were deliberately chosen to test our agent’s ability to adapt to new, unseen tasks. Impressively, our agent successfully completed them with an average success rate (SR) of 95.3%. These results highlight the advantages of our modular architecture. Unlike other approaches that rely on end-to-end imitation learning from human demonstrations, where vision recognition and action planning are tightly integrated, our method divides these functions into distinct modules. This separation allows our agent to leverage the knowledge and reasoning power of the LLM, enabling it to handle unseen tasks without requiring additional training, as long as those tasks don’t introduce entirely new types of UI elements that the visual observer module fails to recognize.

### v3

@@ 1 Introduction
A promising approach to ensure the applicability of agents for more general computer tasks is to rely solely on visual information from the screen as input and use keyboard and mouse actions as output (Shaw et al., 2023; Cheng et al., 2024). By leveraging well-established GUI environments, such agents can be readily deployed to assist humans in performing their computer tasks. Prior research in this direction has primarily focused on using Vision-Transformers, which are a natural choice for processing screenshot images. However, even state-of-the-art Vision-Language Models (VLMs) struggle to ground actions to specific pixel coordinates (AnthropicDocumentation; zheng2024seeact). Without a robust and reliable mechanism for processing the spatial information of GUI elements, achieving a practical agent remains unattainable.
In this paper, we propose a novel agent architecture composed of a separate visual observation module and a reasoning model (LLM), which strongly reinforces the agent’s coordinate-grounding functionality. The visual observation module interprets the environment and represents it in textual form, alleviating the burden on the VLM to simultaneously process visual information and perform task-specific, reasoning-based planning. This modular approach also greatly improves the agent’s adaptability a cross various applications, as each module can be trained or upgraded independently. To further support and maximize the effectiveness of our agent design, we have developed a novel prompting technique called Context-Aware Action Planning (CAAP). This technique systematically organizes the contextual information required for action-proposal and identifies syntactic structures that trigger the most effective Chain of Thought (CoT) (Wei et al., 2022). Our agent demonstrates superior task-solving performance on the MiniWoB++ and WebShop benchmark. It surpasses previous studies by utilizing only a minimal amount of human demonstration data, thereby illustrating its efficiency and effectiveness in handling complex computer tasks with limited training cost.
@@ 2 Related work
Humphreys et al. (2022) introduced CC-Net, a compound model of ResNet, Transformer, and LSTM that uses image and DOM data as inputs and produces policies for keyboard-mouse actions. It was trained via SL and RL, covering the largest number of MiniWoB++ tasks to date (104 tasks). Furuta et al. (2023) proposed WebGUM, which combines Vision Transformer (ViT) and T5 transformers. It was trained using SL to solve both MiniWoB++ and WebShop tasks, utilizing both image and HTML inputs. Shaw et al. (2023) introduced Pix2Act, the first agent to completely eliminate reliance on DOM or HTML for solving MiniWoB++ and WebShop tasks. The Pix2Act model is built upon the Pix2Struct architecture (Lee et al., 2023) and uses only image inputs. While these studies provide valuable insights and serve as important benchmarks, their methods are impractical for real-world applications as they require around a million or more human demonstration data samples, demanding thousands of hours of human labor for training. Cheng et al. (2024) introduced SeeClick, which is aided by the strong reasoning capabilities of the VLM. SeeClick required a much smaller dataset of 2.8K human demonstrations to train for MiniWoB++ tasks. It fell short in overall performance, however, achieving less than a 70% success rate on the 55 tasks it covered. A common feature of all these approaches is the integration of image interpretation and action planning models within a single architecture. This rigid design poses a significant limitation: when new UI elements need to be recognized or additional actions supported for a new task, the entire architecture must be retrained.
@@ 3 Modularized GUI agent for computer tasks
Figure 1: The architecture and task-solving flow of the CAAP agent. The agent interprets a screenshot captured in the computer environment through the visual observer. The action proposer leverages the reasoning capabilities of the LLM to determine the next actions to take based on the observed state. Once actions are decided, the action executer applies the corresponding keyboard and mouse actions to the environment via the OS interface. This sequence of processes across the three modules continues until the task is completed.
Our agent is composed of three distinct modules: the visual observer, which processes and interprets image inputs; the action proposer, which suggests subsequent actions; and the action executor, which carries out the proposed actions. The agent operates by iteratively repeating a cycle in which these modules execute sequentially. Each module is further divided into sub-modules based on its functionality, as illustrated in Figure 1. This modular architecture provides substantial benefits over monolithic agent designs, such as those utilizing VLMs, particularly in flexibility, resource efficiency, and scalability. Modular designs enable targeted updates to specific components, ensuring more robust and stable improvements without impacting the entire system. Additionally, they reduce computational complexity during training. When needed, functional sub-modules can be seamlessly integrated alongside existing ones to further enhance system performance.
@@
The visual observer module employs a two-step approach to transform visual information from a screen into the linguistic domain. Initially, in the detection stage, UI elements are identified, and their spatial locations are extracted. Subsequently, these elements undergo further analysis through advanced object understanding techniques, during which detailed attributes of each element are captured. This processing results in the information being organized into structured text, ready to be handed to the action proposer as the representation of the current state.
**footnotetext: While VLMs may eventually advance enough to replace the visual observer in the future, it would still serve as a valuable complementary tool for interpreting screenshots. For instance, the action proposer’s LLM model could be replaced with a VLM capable of processing both the screenshot image and the textual representation provided by the visual observer, achieving a synergistic integration of multimodal capabilities.
@@
(a) An example of the masking style of the original Pix2Struct that covers the target element with an X-box.
(b) Our semi-masking approach that highlights the target element by outlining it and darkening its surrounding.
(c) An example of the extracted features from the image of (b) by our visual observer.
Figure 2: Comparison of the masking methods for the original Pix2Struct and our UI element understanding model, and an example of the extracted features for our vision observer. While the original Pix2Struct outputs text in a HTML-like format, our model is finetuned to return JSON-style text.
@@
We use the Pix2Struct model (Lee et al., 2023) to extract detailed attributes from all UI elements identified during the detection stage. We fine-tune the original Pix2Struct model to match the specific output format required for our task. During this fine-tuning process, we employ a different masking technique from Pix2Struct style, as depicted in Figure 2 (a). Our masking technique highlights the target UI element by outlining its region and darkening the surrounding image areas, as demonstrated in Figure 2 (b). This method preserves the pixel-level details of the targeted element while helping the model learn the spatial relationships between the element and its adjacent elements. The structured output from the visual observer is presented as shown in Figure 2 (c), with the range of extractable attributes for each element detailed in Table 4 of Appendix A.
3.2 Action proposer
Upon receiving the representation of the current screen, the action proposer is employed to determine possible actions for completing the task. The proposal process is facilitated by its two key components: a context collector and a CAAP prompter. The context collector gathers vital contextual information that aids in task resolution. This includes exemplary demonstrations, the ultimate task goal, the history of actions performed so far, and the range of possible actions that the language model can choose from, all formatted as text. The CAAP prompter then uses this information to construct prompts that are sent to the language model, eliciting responses that guide the next steps. The CAAP prompting method, one of our major contributions in this paper, is discussed in detail in Section 4.
@@ 3.3 Action executor
The actions determined by the action proposer are implemented through a low-level interface to effect changes in the environment by the last module, action executor. It first identifies a series of recommended actions by parsing the response text from the action proposer. Then, it executes these actions using the keyboard and mouse interfaces of the operating system. After each action, it receives an updated screenshot, halting further actions if a change in the environment is detected. The types of actions that can be executed include mouse actions such as clicking, pointing, and scrolling, as well as keyboard actions like typing and using common shortcuts. These are enumerated in Table 5 of Appendix B and implemented as Python functions using the PyAutoGUI package.
@@ 4 CAAP prompting
Figure 3: Content design for the CAAP prompting.
The performance of LLM-based agents is significantly influenced by the structure of their prompts. Our prompting approach is specifically designed to enable the LLM to effectively utilize surrounding context information essential for decision-making. This approach aims to elicit a CoT reasoning process. The strength of the CoT mechanism lies in its ability to connect disparate pieces of context information, progressively reinforcing their cohesion and enriching the reasoning process. CAAP prompting is specifically developed to enable the effective utilization of this mechanism during the action proposal phase. To achieve this, it actively incorporates both explicit and implicit strategies to guide the LLM’s CoT reasoning. The design of CAAP prompting is systematically structured as illustrated in Figure 3.
4.1 Explicit and implicit CoT inducement
Explicit inducement of CoT
Inspired by human action planning, we identified four essential types of surrounding context information necessary for determining subsequent actions. These contextual elements are conveyed to the LLM along with CoT-inducing instructions, which explicitly guide the model to establish semantic connections between them. Using CoT-inducing instructions, the LLM examines the actions taken so far to uncover the relationship between the task objective and the action trajectory. It then infers the contextual meaning of these actions and the updated state of the visual environment. Subsequently, referencing the candidates of action types, it refines the subsequent action plan considering the current situation and outputs actionable guidance in a structured text format.
Implicit inducement of CoT
In addition to CoT-inducing instructions, we enhance the implicit induction of chain-of-thought (CoT) reasoning in large language models (LLMs) through human demonstrations. By providing exemplars of few-shot task-solving action trajectories from experts, LLMs can be guided to reason step-by-step Wei et al. (2022). However, an over-reliance on demonstrations can lead LLMs to rigidly mimic the examples, making them less capable of adapting to exceptional scenarios. To address this, we augment each action in the demonstrations with corresponding rationales, enabling the LLM to better comprehend the contextual information relevant to the action. These rationales contain detailed natural language descriptions of the actions and the associated circumstances. The methodology for generating these rationales using LLMs is described in the ’human demo dataset’ section of Appendix C.
The prompts both explicitly and implicitly inducing CoT reasoning and the subsequent inference processes of the LLM can be found in Appendix D, specifically in Figures 6 and 7, respectively.
4.2 CAAP components
@@
Providing few-shot exemplars within the prompt is a common practice that enhances the advanced reasoning capabilities of LLMs. CAAP prompt begins with few-shot human demonstrations, which are sequences of actions with rationales for solving tasks similar to the given objective.
@@
This part contains all the essential information for the language model to propose the next actions. It includes a description of the target task, a record of past actions, state descriptions from the visual observer, and the available range of actions for the agent.
@@
The section on CoT-inducing instructions contains phrases that encourage the LLM to link the surrounding context information, thereby inducing CoT. Instruction for reviewing action trajectory encourages the agent to review its actions taken so far. Instruction for reviewing visual observation prompts the agent to examine the current state’s observations. Instruction for improving action plan requests the formulation of a subsequent action plan, taking into account the self-generated text thus far. Finally, the Instruction for deciding next actions provides guidance on the desired response format and requests a list of actions to be taken in the subsequent steps.
@@ 5.1 MiniWoB++
MiniWoB++ is a highly suitable benchmark for evaluating an agent’s ability to handle a wide range of scenarios encountered during computer task execution, as it encompasses over a hundred of fundamental tasks (Liu et al., 2018; Shi et al., 2017; Min, ). As an extension of OpenAI MiniWoB, this benchmark includes a diverse spectrum of tasks, ranging from simple activities such as button clicks and form filling to more complex tasks like email forwarding and flight booking. We utilize the MiniWoB++ benchmark to thoroughly assess our agent’s effectiveness in navigating and solving diverse tasks within a simulated computer environment.
Previous studies using the MiniWoB++ benchmark often choose subsets of tasks most suitable for demonstrating their methods. We have also selected a total of 73 MiniWoB++ tasks that meet the following criteria: (1) tasks that consist of UI elements of types recognizable by our visual observer module, as listed in Table 4 of Appendix A (2) tasks where none of UI elements extend beyond the default screen area of 160x210 pixels (3) tasks that do not require color recognition, as our visual observer module is not yet trained for this capability.
@@
For the action-proposal language model, we utilize the Azure OpenAI gpt-4-0125 model. The Azure OpenAI API features the ‘function calling’ capability, thus our prompt does not directly include the list of action types (as depicted in Figure 3). Instead, this contextual information is conveyed to the language model through arguments in the function calling feature. An example of how the function calling is applied is illustrated in Figure 8 in Appendix E.
@@
Table 1 compares the performance of our approach with other studies reported on the MiniWoB++ benchmark. Since each study reports results over different sets of tasks (See Figure 9 of Appendix F.), we have included both the SRs reported in the literature and the number of tasks used in each case. First, CC-Net, which covers the largest set of tasks, was trained on 2.4M human demonstration data across 104 tasks. However, when using only image input (excluding DOM), its performance dropped to 24%, indirectly highlighting the challenge of solving computer tasks using image inputs alone. Pix2Act, trained on 1.3M human demonstrations across 59 tasks, achieved an SR of 96.2%, demonstrating the feasibility of agents solving computer tasks with image input only. SeeClick, despite using a relatively small dataset of 2.8K demonstrations to learn click locations on screenshots, achieved an SR of only 69.4% across 55 tasks, showing limited effectiveness. In contrast, our approach, using only 97 demonstration samples, achieved a high task SR of 94.5% across 73 tasks. Although an additional 1.76K samples***These image samples were semi-automatically annotated using a custom tool shown in Figure 10 of Appendix H. were required to train the model for image input interpretation due to the modular nature of our approach, our method still demonstrated its effectiveness on a broader range of computer tasks with about 1.86K samples in total. Task-level SRs are detailed in Table 7 of Appendix G.
Of the 73 tasks included in our experiments, six tasks—drag-circle, drag-single-shape, find-greatest, generate-number, odd-or-even, and sign-agreement— were excluded during the generation of training data for the visual observer, making them unseen tasks for our agent. Despite this, the agent successfully completed these tasks with an average success rate of 95.3%. These results underscore the benefits of our LLM-based approach compared to the prior end-to-end learning method that relied on imitation learning from human demonstrations. As long as the new tasks do not introduce completely unfamiliar types of UI elements that the visual observer cannot recognize, our agent can effectively manage these unseen tasks by utilizing the knowledge and reasoning capabilities of the LLM.